OpenStudy (anonymous):

Solve the equation: 7 - y - y = -1

OpenStudy (cj49):

7-2y=-1 -2y=-1-7 -2y=-8 y=8/2=?

Nnesha (nnesha):

combine like terms which is -y and -y mean -1-1 =??